OpenStudy (anonymous):

(3,18) is this an ordered pair of 3y-4x>6

OpenStudy (anonymous):

yes..... i think

myininaya (myininaya):

x=3 y=18 plug these numbers in and see if it is greater than 6

myininaya (myininaya):

if it is greater than 6 then yes if it is not then no

OpenStudy (anonymous):

3*18 - 4*3 = 54 - 12 54 - 12 = 42 42 > 6